# Agentic Retoucher for Text-To-Image Generation

### v1

@@ 1 Introduction
Text-to-image (T2I) diffusion models such as Imagen [saharia2022photorealistic], Stable Diffusion [rombach2022high, podell2023sdxl], FLUX [flux2024], and Qwen-Image [wu2025qwenimagetechnicalreport] have revolutionized image synthesis, enabling photorealistic and creative generation from natural language prompts. They are now widely adopted in design, film, and entertainment pipelines, as well as in downstream tasks like editing [brooks2023instructpix2pix, kawar2023imagic, deng2025bagel, liu2025step1x-edit] and video generation [wan2025, opensora, chen2025sanavideoefficientvideogeneration]. However, even the most advanced models frequently produce small-scale distortions, including misaligned limbs, asymmetric faces, unreadable text, and inconsistent object interactions. These flaws typically occur locally within otherwise high-quality outputs, making them difficult to detect and expensive to correct through full-image regeneration. As a result, T2I systems still lack autonomous perceptual reliability, a key barrier to real-world creative and industrial use.
Recent research has explored three main directions to improve generative fidelity: prompt enhancement [hertz2022prompt, wang2025promptenhancersimpleapproachenhance, wu2025reprompt], reinforcement learning-based optimization [black2023ddpo], and fine-grained noise-space alignment [izadi2025finegrainedalignmentnoiserefinement, Le2025FromRT, wu2024megafusion]. Although these approaches effectively enhance overall realism, they lack explicit spatial reasoning and cannot interpret or correct localized failures. Post-hoc editing pipelines such as Imagic [kawar2023imagic], Bagel [deng2025bagel], and Step1x-Edit [liu2025step1x-edit] enable local refinement, but rely on manually crafted masks or heuristic textual hints, preventing autonomous identification of regions requiring correction.
Vision-language models (VLMs) [openai2024gpt4ocard, li2024llavaonevisioneasyvisualtask] show promise as automated critics due to their semantic reasoning capabilities. However, as shown in Fig. 1 (Left), even state-of-the-art VLMs struggle to reliably localize distorted regions. Explicit queries often yield inconsistent or incorrect assessments, with clearly abnormal regions being misjudged as normal. This stems from two key issues: VLMs are optimized for high-level semantic alignment rather than pixel-level verification, leading to weak spatial grounding and missed fine-scale artifacts. Furthermore, their extensive knowledge priors can override visual evidence, causing hallucinated judgments. For example, a portrait with six fingers is deemed plausible despite explicit highlighting of the defective hand, demonstrating that current VLMs are unreliable for fine-grained artifact detection in AI-generated images.
@@ 2 Related Works
Visual Quality Assessment. Visual Quality Assessment (VQA) [qian2025explainablepartialaigcimagequality, Liang_2024_CVPR, cao2024synartifact, chang2025oneigbenchomnidimensionalnuancedevaluation] is an important and rapidly evolving field that has made significant contributions to evaluating a wide range of image and video tasks with closer alignment to human subjective perception. For AIGC content assessment, most existing work [wang2025cigeval, zhang2024abench] is limited to applying quantitative metrics at global scales, without explicit localization and assessment of local flaws. RichHF [Liang_2024_CVPR] introduces predictors for local structural distortions along with a corresponding scoring procedure. However, these methods focus solely on assessment and have not been integrated into an automated, closed-loop pipeline for evaluation and refinement.
Vision-Language Model (VLMs). VLMs [vteam2025glm45vglm41vthinkingversatilemultimodal, yang2025qwen3technicalreport, lu2025ovis25technicalreport] have become leading drivers of general artificial intelligence, exhibiting remarkable problem-solving and reasoning ability through training on large-scale multimodal data (e.g., GPT-4o [GPT4] and the Qwen [yang2025qwen3technicalreport] family in real-world multimodal interaction). Several works [kaduri2024_vision_of_vlms, Qu_2025_CVPR, Heo_2025_CVPR] further advance VLMs on image understanding tasks. However, heavy reliance on high-fidelity pretraining data and learned priors often biases VLMs toward prior-based, ungrounded hallucinated responses in the context of text-to-image evaluation.
Agentic System in Vision. Agentic System [538972, RotherKB04, Cubuk2019AutoAugmentLA, wang2021tent, zhu2023ghost, li2023generalist, chen2024restoreagent] adopts active, closed-loop perception-decision-action framework, with VLMs increasingly acting as planners due to their strong reasoning. In the 3D domain, VADAR [Marsili_2025_CVPR] proposes an agentic program synthesis approach, achieving superior performance in 3D spatial reasoning. In image and video restoration, AgenticIR [agenticir] and MoA-VR [liu2025moavrmixtureofagentsallinonevideo] independently propose VLM-integrated multi-agent repair paradigms. In the realm of artistic creation, JarvisArt [jarvisart2025] enables fine-grained photo retouching via tool invocation based on user instructions.
@@ 3.1 Distortion Taxonomy
Existing T2I evaluation datasets suffer from limited coverage (e.g., HADM [Wang2024HADM], Wang et al. [Wang_2025_CVPR]), coarse annotation (e.g., RichHF [Liang_2024_CVPR]), and insufficient scale (e.g., SynArtifacts-1K [cao2024synartifact] with only 1K samples). To address these issues, GenBlemish-27K establishes a hierarchical taxonomy of distortions derived from large-scale inspection of outputs from mainstream T2I models. We define six high-level distortion dimensions including human anatomical distortion, attribute inconsistency, spatial errors, object deformation or redundancy, action and interaction distortion, and miscellaneous cases. These dimensions are further refined into 12 fine-grained categories such as limb deformities, face distortion, and text anomalies. This taxonomy captures typical artifacts observed in state-of-the-art diffusion models and enables interpretable reasoning about both what and where fails (see Fig. 2).
@@ 3.2 Data Annotation
We curate 6,025 images from EvalMuse-Structure [han2025ntire2025challengetext], covering outputs from over 20 T2I models such as Dreamina, Midjourney, Kandinsky [arkhipkin2023kandinsky], and SDXL [podell2023sdxl]. A four-stage human-in-the-loop process ensures both semantic richness and annotation consistency. (1) Annotators are first calibrated through a pre-annotation stage. (2) For each distortion region, multiple annotators independently provide the center, category, and a brief textual description; the region radius is 1/20 of the image height. (3) The textual descriptions are expanded and refined using QwenVL-Max [Qwen2.5-VL]. (4) Final annotations are reconciled through majority voting and expert validation. Each sample includes the generated image, input prompt, distortion mask, category label, and natural-language description, supporting tasks such as saliency prediction, defect classification, and language grounding.
@@ 4.2 Context-Aware Perceptual Distortion Analysis
Text-to-image generations frequently exhibit subtle and context-dependent distortions such as implausible limb, object and text. These artifacts often lack explicit object boundaries, making conventional pixel-wise detection unreliable. To emulate human perceptual sensitivity, inspired by [Liang_2024_CVPR], we design a context-aware saliency predictor that estimates a distortion-saliency map $S\in[0,1]^{H\times W}$ conditioned on both the generated image $I$ and its prompt $P$. A dual-encoder ViT [2020An]-T5 [roberts2022scalingmodelsdatatextttt5x] backbone encodes image and text representations, which are subsequently concatenated and fused via a self-attention mechanism to capture inherent correspondences between visual structures and textual semantics. A lightweight attention refinement module further aggregates multi-scale contextual cues, improving the detection of distortions whose visibility depends on global images.
@@ 4.3 Human-Aligned Reasoning and Adaptive Action
In the first stage, SFT establishes structured response formats and distortion taxonomy under limited supervision. To reduce computational overhead, we employ Low-Rank Adaptation (LoRA) [hu2022lowrank], where the weight update $\Delta W$ for a layer $W\in\mathbb{R}^{n\times m}$ is decomposed as $\Delta W=AB$, with $A\in\mathbb{R}^{n\times r}$ and $B\in\mathbb{R}^{r\times m}$. This low-rank decomposition enables efficient specialization of the reasoning model without full-parameter fine-tuning.
In the second stage, GRPO [deepseekai2025deepseekr1incentivizingreasoningcapability] aligns reasoning behavior with human preferences through reinforcement signals:
@@ 5.1 Experimental setup
Datasets and Implementation Details. We evaluate our framework on the proposed GenBlemish-27K dataset to assess optimization efficacy and module functionality, and further verify its generalization on SynArtifacts-1K [cao2024synartifact]. For the Context-Aware Perception Agent, the learning rate is set to $2\times 10^{-5}$. For the Human-Alignment Reasoning Agent, Stage 1 employs LoRA fine-tuning with rank 64 and $\alpha=32$. Inference is fully automatic and converges within 2-3 reasoning iterations per image.
Evaluation Metrics. For retouch evaluation, we use the four perceptual metrics from RichHF [Liang_2024_CVPR]: plausibility, aesthetics, alignment, and overall, to assess both structural plausibility and perceptual quality, as standard T2I metrics (e.g., FID [NIPS2017_8a1d6947]) fail to capture localized improvements. For perception, following [8315047], we adopt CC, SIM, KLD, AUC-Judd, and NSS to evaluate distributional and fixation-level consistency. For reasoning, we report distortion-type classification accuracy and semantic alignment using ROUGE [lin-2004-rouge], METEOR [banerjee-lavie-2005-meteor], Word2Vec [mikolov2013efficientestimationwordrepresentations], and SimCSE [gao2021simcse]. More setup details are provided in the supplementary material.
@@ 5.2 Comparison
SALICON [Huang_2015_ICCV] 0.9230 1.0774 0.5039 0.2734 1.7171
TranSalNet [Lou2022TranSalNetTP] 0.9029 1.1494 0.4616 0.0989 2.8716
Sal-CFS-GAN [Che_2020] 0.7747 0.7810 0.2124 0.1018 2.8589
SAM-VGG [cornia2018predicting] 0.8773 1.2072 0.3410 0.1791 2.4094
SAM-ResNet [cornia2018predicting] 0.9162 1.0552 0.4040 0.2475 2.0740
MLNet [mlnet2016] 0.8539 1.0455 0.3535 0.2381 2.2359
InternVL3.5-8B [wang2025internvl35advancingopensourcemultimodal] 0.8049 0.7689 0.5104 0.4095 3.9325
Qwen2.5-VL-7B [Qwen2.5-VL] 0.6145 0.4190 0.1710 0.1481 7.4353
GLM-4.1V-9B [vteam2025glm45vglm41vthinkingversatilemultimodal] 0.5461 0.2191 0.0604 0.0902 8.0118
RichHF [Liang_2024_CVPR] 0.9211 0.8954 0.4748 0.3309 1.6697
@@ 5.3 Perception and Reasoning Analysis
Context-Aware Perception. Tab. 3 compares the proposed Context-Aware Perception Agent with conventional saliency detectors, deep saliency networks, and vision-language models. Hand-crafted methods (e.g., AIM [NIPS2005_0738069b], GBVS [NIPS2006_4db0f8b0]) rely on low-level contrast, while deep saliency networks (e.g., SAM-ResNet [cornia2018predicting], TranSalNet [Lou2022TranSalNetTP]) yield only moderate gains; both fail to capture context-aware distortion regions. General-purpose VLMs (e.g., InternVL3.5 [wang2025internvl35advancingopensourcemultimodal], Qwen2.5-VL [Qwen2.5-VL], GLM-4.1V [vteam2025glm45vglm41vthinkingversatilemultimodal]) also perform poorly, lacking task-specific grounding for distortion awareness. In contrast, our perception agent achieves state-of-the-art results across all metrics (AUC-Judd = 0.9336, NSS = 1.2087, KLD = 1.4313), demonstrating robust localization of artifact-prone regions.
@@
Gemini-2.5 Pro Zero-Shot [comanici2025gemini25pushingfrontier] 60.28% 0.6856 0.6245 0.1702 0.1121
Qwen2.5-VL-7B [Qwen2.5-VL] 57.76% 0.6658 0.6110 0.1678 0.0733

### v3

@@ Abstract
00footnotetext: *: These authors contribute equally to this work.
$\dagger$: Corresponding author.
@@ 1 Introduction
Text-to-image (T2I) diffusion models [22, 41, 23] such as Imagen [61], Stable Diffusion [59, 54], FLUX [37], and Qwen-Image [57] have revolutionized image synthesis, enabling photorealistic and creative generation from natural language prompts. They are now widely adopted in design, film, and entertainment pipelines, as well as in downstream tasks like editing [5, 36, 19, 47, 74] and video generation [64, 77, 13]. However, even the most advanced models frequently produce small-scale distortions, including misaligned limbs, asymmetric faces, unreadable text, and inconsistent object interactions. These flaws typically occur locally within otherwise high-quality outputs, making them difficult to detect and expensive to correct through full-image regeneration. As a result, T2I systems still lack autonomous perceptual reliability, a key barrier to real-world creative and industrial use.
Recent research has explored three main directions to improve generative fidelity: prompt enhancement [28, 68, 72], reinforcement learning-based optimization [4], and fine-grained noise-space alignment [34, 38, 71]. Although these approaches effectively enhance overall realism, they lack explicit spatial reasoning and cannot interpret or correct localized failures. Post-hoc editing pipelines such as Imagic [36], Bagel [19], and Step1x-Edit [47] enable local refinement, but rely on manually crafted masks or heuristic textual hints, preventing autonomous identification of regions requiring correction.
Vision-language models (VLMs) [53, 39] show promise as automated critics due to their semantic reasoning capabilities. However, as shown in Fig. 1 (Left), even state-of-the-art VLMs struggle to reliably localize distorted regions. Explicit queries often yield inconsistent or incorrect assessments, with clearly abnormal regions being misjudged as normal. This stems from two key issues: VLMs are optimized for high-level semantic alignment rather than pixel-level verification, leading to weak spatial grounding and missed fine-scale artifacts. Furthermore, their extensive knowledge priors can override visual evidence, causing hallucinated judgments. For example, a portrait with six fingers is deemed plausible despite explicit highlighting of the defective hand, demonstrating that current VLMs are unreliable for fine-grained artifact detection in AI-generated images.
@@ 2 Related Works
Visual Quality Assessment. Visual Quality Assessment (VQA) [55, 42, 8, 10, 21, 46] is an important and rapidly evolving field that has made significant contributions to evaluating a wide range of image and video tasks with closer alignment to human subjective perception. For AIGC content assessment, most existing work [66, 75] is limited to applying quantitative metrics at global scales, without explicit localization and assessment of local flaws. RichHF [42] introduces predictors for local structural distortions along with a corresponding scoring procedure. However, these methods focus solely on assessment and have not been integrated into an automated, closed-loop pipeline for evaluation and refinement.
Vision-Language Model (VLMs). VLMs [63, 73, 49] have become leading drivers of general artificial intelligence, exhibiting remarkable problem-solving and reasoning ability through training on large-scale multimodal data (e.g., GPT-4o [62] and the Qwen [73] family in real-world multimodal interaction). Several works [35, 56, 27] further advance VLMs on image understanding tasks. However, heavy reliance on high-fidelity pretraining data and learned priors often biases VLMs toward prior-based, ungrounded hallucinated responses in the context of text-to-image evaluation.
Agentic System in Vision. Agentic System [33, 60, 17, 65, 79, 40, 12] adopts active, closed-loop perception-decision-action framework, with VLMs increasingly acting as planners due to their strong reasoning. In the 3D domain, VADAR [50] proposes an agentic program synthesis approach, achieving superior performance in 3D spatial reasoning. In image and video restoration, AgenticIR [78] and MoA-VR [45] independently propose VLM-integrated multi-agent repair paradigms. In the realm of artistic creation, JarvisArt [44] enables fine-grained photo retouching via tool invocation based on user instructions.
@@ 3.1 Distortion Taxonomy
Existing T2I evaluation datasets suffer from limited coverage (e.g., HADM [67], Wang et al. [70]), coarse annotation (e.g., RichHF [42]), and insufficient scale (e.g., SynArtifacts-1K [8] with only 1K samples). To address these issues, GenBlemish-27K establishes a hierarchical taxonomy of distortions derived from large-scale inspection of outputs from mainstream T2I models. We define six high-level distortion dimensions including human anatomical distortion, attribute inconsistency, spatial errors, object deformation or redundancy, action and interaction distortion, and miscellaneous cases. These dimensions are further refined into 12 fine-grained categories such as limb deformities, face distortion, and text anomalies. This taxonomy captures typical artifacts observed in state-of-the-art diffusion models and enables interpretable reasoning about both what and where fails (see Fig. 2).
@@ 3.2 Data Annotation
We curate 6,025 images from EvalMuse-Structure [52], covering outputs from over 20 T2I models such as Dreamina, Midjourney, Kandinsky [1], and SDXL [54]. A four-stage human-in-the-loop process ensures both semantic richness and annotation consistency. (1) Annotators are first calibrated through a pre-annotation stage. (2) For each distortion region, multiple annotators independently provide the center, category, and a brief textual description; the region radius is 1/20 of the image height. (3) The textual descriptions are expanded and refined using QwenVL-Max [2]. (4) Final annotations are reconciled through majority voting and expert validation. Each sample includes the generated image, input prompt, distortion mask, category label, and natural-language description, supporting tasks such as saliency prediction, defect classification, and language grounding.
@@ 4.2 Context-Aware Perceptual Distortion Analysis
Text-to-image generations frequently exhibit subtle and context-dependent distortions such as implausible limb, object and text. These artifacts often lack explicit object boundaries, making conventional pixel-wise detection unreliable. To emulate human perceptual sensitivity, inspired by [42], we design a context-aware saliency predictor that estimates a distortion-saliency map $S\in[0,1]^{H\times W}$ conditioned on both the generated image $I$ and its prompt $P$. A dual-encoder ViT [20]-T5 [58] backbone encodes image and text representations, which are subsequently concatenated and fused via a self-attention mechanism to capture inherent correspondences between visual structures and textual semantics. A lightweight attention refinement module further aggregates multi-scale contextual cues, improving the detection of distortions whose visibility depends on global images.
@@ 4.3 Human-Aligned Reasoning and Adaptive Action
In the first stage, SFT establishes structured response formats and distortion taxonomy under limited supervision. To reduce computational overhead, we employ Low-Rank Adaptation (LoRA) [31], where the weight update $\Delta W$ for a layer $W\in\mathbb{R}^{n\times m}$ is decomposed as $\Delta W=AB$, with $A\in\mathbb{R}^{n\times r}$ and $B\in\mathbb{R}^{r\times m}$. This low-rank decomposition enables efficient specialization of the reasoning model without full-parameter fine-tuning.
In the second stage, GRPO [18] aligns reasoning behavior with human preferences through reinforcement signals:
@@ 5.1 Experimental setup
Datasets and Implementation Details. We evaluate our framework on the proposed GenBlemish-27K dataset to assess optimization efficacy and module functionality, and further verify its generalization on SynArtifacts-1K [8]. For the Context-Aware Perception Agent, the learning rate is set to $2\times 10^{-5}$. For the Human-Alignment Reasoning Agent, Stage 1 employs LoRA fine-tuning with rank 64 and $\alpha=32$. Inference is fully automatic and converges within 2-3 reasoning iterations per image.
Evaluation Metrics. For retouch evaluation, we use the four perceptual metrics from RichHF [42]: plausibility, aesthetics, alignment, and overall, to assess both structural plausibility and perceptual quality, as standard T2I metrics (e.g., FID [29]) fail to capture localized improvements. For perception, following [7], we adopt CC, SIM, KLD, AUC-Judd, and NSS to evaluate distributional and fixation-level consistency. For reasoning, we report distortion-type classification accuracy and semantic alignment using ROUGE [43], METEOR [3], Word2Vec [51], and SimCSE [24]. More setup details are provided in the supplementary material.
@@ 5.2 Comparison
SALICON [32] 0.9230 1.0774 0.5039 0.2734 1.7171
TranSalNet [48] 0.9029 1.1494 0.4616 0.0989 2.8716
Sal-CFS-GAN [11] 0.7747 0.7810 0.2124 0.1018 2.8589
SAM-VGG [16] 0.8773 1.2072 0.3410 0.1791 2.4094
SAM-ResNet [16] 0.9162 1.0552 0.4040 0.2475 2.0740
MLNet [15] 0.8539 1.0455 0.3535 0.2381 2.2359
InternVL3.5-8B [69] 0.8049 0.7689 0.5104 0.4095 3.9325
Qwen2.5-VL-7B [2] 0.6145 0.4190 0.1710 0.1481 7.4353
GLM-4.1V-9B [63] 0.5461 0.2191 0.0604 0.0902 8.0118
RichHF [42] 0.9211 0.8954 0.4748 0.3309 1.6697
@@ 5.3 Perception and Reasoning Analysis
Context-Aware Perception. Tab. 3 compares the proposed Context-Aware Perception Agent with conventional saliency detectors, deep saliency networks, and vision-language models. Hand-crafted methods (e.g., AIM [6], GBVS [26]) rely on low-level contrast, while deep saliency networks (e.g., SAM-ResNet [16], TranSalNet [48]) yield only moderate gains; both fail to capture context-aware distortion regions. General-purpose VLMs (e.g., InternVL3.5 [69], Qwen2.5-VL [2], GLM-4.1V [63]) also perform poorly, lacking task-specific grounding for distortion awareness. In contrast, our perception agent achieves state-of-the-art results across all metrics (AUC-Judd = 0.9336, NSS = 1.2087, KLD = 1.4313), demonstrating robust localization of artifact-prone regions.
@@
Gemini-2.5 Pro Zero-Shot [14] 60.28% 0.6856 0.6245 0.1702 0.1121
Qwen2.5-VL-7B [2] 57.76% 0.6658 0.6110 0.1678 0.0733
@@ 6 Conclusions
Acknowledgment
This work is supported by National Natural Science Foundation of China (62571322, 62431015, 62271308), STCSM (24ZR1432000, 24511106902, 24511106900, 22DZ2229005), 111 plan (BP0719010), and State Key Laboratory of UHD Video and Audio Production and Presentation.